OpenStudy (anonymous):

FAN AND MEDAL!!!!!! Pls help What is the midpoint between (8, -5) and (2, -2)? (5, -1.5) (10, -7) (3, -1.5) (5, -3.5)

OpenStudy (alexandervonhumboldt2):

midpoint formula ((x1+x2)/2 , (y1+y2)/2)

OpenStudy (igreen):

Use the formula \((\dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2})\)

OpenStudy (igreen):

Plug in what we know: \((\dfrac{8 + 2}{2}, \dfrac{-5 + -2}{2})\) Simplify

OpenStudy (igreen):

Can you simplify that? @Bandmom_4123

OpenStudy (anonymous):

Yea I gt it. Thnx

OpenStudy (igreen):

No problem.